OpenStudy (anonymous):

Write the equation of the quadratic function with roots -9 and and -3 and a vertex at (-6, -1). Answer

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

you can always write parabola as \[y=a \left( x-x _{1} \right)\left( x-x _{2} \right)\] where x1 and x2 are roots of parabola and a is just constant

OpenStudy (anonymous):

so it would be y=a(x+9)(x+3)

OpenStudy (anonymous):

yes thats right and now substitute vortex point in the ecuation and you will get a

OpenStudy (anonymous):

-6(x+9)(x+3) like tht?

OpenStudy (anonymous):

did you substitute -6 for x and -1 for y? because i got different a

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

i got f(x)=x^2-12x+27

OpenStudy (anonymous):

a(-6+9)(-6+3)=-1

OpenStudy (anonymous):

a(3*-3)=-1 a(-9)=-1

OpenStudy (anonymous):

a=8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why are you adding 9 when you should divide by 9

OpenStudy (anonymous):

so its 1/9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks now wat do i do if i want the full equation

OpenStudy (anonymous):

multiply out (1/9)(x+9)(x+6)

OpenStudy (anonymous):

ok just a moment

OpenStudy (anonymous):

not x+6 a should write x+3

OpenStudy (anonymous):

sorry :)

OpenStudy (anonymous):

ok i got

OpenStudy (anonymous):

x^2/9+4x+3+3

OpenStudy (anonymous):

i wrote wrong equation up (1/9)(x+3)(x+9) :D sorry about that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im so confused can you tell me the equation you got

OpenStudy (anonymous):

what confusis you? y=a(x-x1)(x-x2) you know a x1 and x2 put them in you will get y=(1/9)(x+3)(x+9) now multiply it out and you will get answer

OpenStudy (anonymous):

y=1/9x^2-1.3x+3

OpenStudy (anonymous):

4/3 is not 1.3

OpenStudy (anonymous):

i think you have some problems in arithamtics

OpenStudy (anonymous):

ok first just multipla (x+9)(x+3)=x^2+3x+9x+27=x^2+12x+3

OpenStudy (anonymous):

This is wrong because i put this in a calculator that determines the equations for quadratic functions and it said that this equation is with the toots of 9,3 and the vertex is -6,1

OpenStudy (anonymous):

which equation y=(1/9)(x+9)(x+3) is wright

OpenStudy (anonymous):

you just multiplied it out wrongly

OpenStudy (anonymous):

you must get x^2/9+4x/3+3

OpenStudy (anonymous):

oh thank u so much got the answer i appreciate your patience

OpenStudy (anonymous):

it was my foult i'm not great at explaning :)